### v1

# Model-free Rayleigh weight from x-ray Thomson scattering measurements

Abstract
X-ray Thomson scattering (XRTS) has emerged as a powerful tool for the diagnostics of matter under extreme conditions. In principle, it gives one access to important system parameters such as the temperature, density, and ionization state, but the interpretation of the measured XRTS intensity usually relies on theoretical models and approximations. In this work, we show that it is possible to extract the Rayleigh weight—a key property that describes the electronic localization around the ions—directly from the experimental data without the need for any model calculations or simulations. As a practical application, we consider an experimental measurement of strongly compressed Be at the National Ignition Facility (NIF) [Döppner et al., Nature 618, 270-275 (2023)]. In addition to being interesting in their own right, our results will open up new avenues for diagnostics from ab initio simulations, help to further constrain existing chemical models, and constitute a rigorous benchmark for theory and simulations.
The x-ray Thomson scattering (XRTS) method Glenzer and Redmer (2009); Sheffield et al. (2010) constitutes a powerful experimental technique, which is capable of giving microscopic insights into a probed sample. A particularly important use case for XRTS is the diagnostics of experiments with matter under extreme densities, temperatures and pressures Falk (2018). Such warm dense matter (WDM) Graziani et al. (2014); Dornheim et al. (2023a) naturally occurs in various astrophysical objects such as giant planet interiors Benuzzi-Mounaix et al. (2014); Guillot et al. (2022); Drake (2018), brown dwarfs Becker et al. (2014) and the outer crust of neutron stars Gudmundsson et al. (1983). Moreover, WDM plays an important role in a variety of technological applications such as inertial confinement fusion (ICF) Betti and Hurricane (2016), where the fuel capsule has to traverse the WDM regime in a controlled way to reach ignition Hu et al. (2011). The recent breakthrough at the National Ignition Facility (NIF) to reach ignition The Indirect Drive ICF Collaboration (2024) has further substantiated the importance of accurately diagnosing WDM states. In the laboratory, these extreme states can be realized using a variety of experimental techniques Falk (2018), and XRTS is often used to infer a-priori unknown system parameters such as the mass density $\rho$, temperature $T$, and ionization state $Z$; see, e.g., Ref. Plagemann et al. (2015). These properties can then be used for physical considerations, to inform equation-of-state tables Falk et al. (2012, 2014); Gaffney et al. (2018), and to benchmark integrated multi-scale simulations such as radiation hydrodynamics Pomraning (2005).
In practice, the measured XRTS intensity can be accurately expressed as Gawne et al. (2024a)
$I(\mathbf{q},E)=S_{ee}(\mathbf{q},E)\circledast R(E)\ ,$ (1)
where $S_{ee}(\mathbf{q},E)$ denotes the electronic dynamic structure factor (DSF) that describes the probed system, and $R(E)$ the combined source-and-instrument function (SIF) that takes into account the shape of the x-ray source and effects of the detector Gawne et al. (2024a). We note that a deconvolution solving Eq. (1) for $S_{ee}(\mathbf{q},E)$ constitutes a major challenge considering the uncertainties of the SIF and the experimental error bars Höll et al. (2007); Fortmann et al. (2009). Therefore, the usual way to interpret the measured XRTS intensity has been to construct a forward model for $S_{ee}(\mathbf{q},E)$, convolve it with $R(E)$, and then fit the resulting trial intensity to the LHS of Eq. (1) where the unknown system properties are treated as free parameters. Naturally, the thus inferred information depends on the employed forward model for $S_{ee}(\mathbf{q},E)$. These are usually based on a number of assumptions such as the possibility to distinguish between bound and free electrons as proposed within the popular Chihara approach Chihara (1987); Höll et al. (2007); Fortmann et al. (2009); Glenzer and Redmer (2009); Kraus et al. (2019); Döppner et al. (2023); Böhme et al. (2023).
Figure 1: Left: XRTS measurement on strongly compressed Be at the NIF Döppner et al. (2023) (green) and its decomposition into elastic (grey) and inelastic (red) contributions. Center: Determination of the normalization $S_{ee}(\mathbf{q})=F_{ee}(\mathbf{q},0)$ from the f-sum rule Dornheim et al. (2024a), cf. Eq. (7). Right: Wavenumber dependence of the Rayleigh weight $W_{R}(\mathbf{q})$, comparing the experimental data point [Eq. (6)] with ab initio PIMC simulations Dornheim et al. (2024b); the solid lines have been obtained for $T=155.5\,$eV and the shaded areas correspond to the uncertainty range of ±$\pm 15\,$eV.
Very recently, it has been proposed to instead consider the two-sided Laplace transform of the DSF Dornheim et al. (2022, 2023b, 2023a),
$F_{ee}(\mathbf{q},\tau)=\mathcal{L}\left[S_{ee}(\mathbf{q},E)\right]=\int_{-\infty}^{\infty}\textnormal{d}E\ S_{ee}(\mathbf{q},E)\ e^{-\tau E}\ ,$ (2)
which is directly related to the imaginary-time density–density correlation function (ITCF) $F_{ee}(\mathbf{q},\tau)$. The latter naturally emerges in Feynman’s path-integral picture of statistical mechanics Kleinert (2009) and, from a physical perspective, contains the same information as $S_{ee}(\mathbf{q},E)$, although in an unfamiliar representation Dornheim et al. (2023c, d). It corresponds to the usual intermediate scattering function $F(\mathbf{q},t)=\braket{\hat{n}(\mathbf{q},t)\hat{n}(-\mathbf{q},0)}$ with an imaginary argument $t=-i\hbar\tau$ with $\tau\in[0,\beta]$ and $\beta=1/k_{\textnormal{B}}T$ the inverse temperature. A key advantage of working in the Laplace domain is given by the convolution theorem
$\mathcal{L}\left[S_{ee}(\mathbf{q},E)\right]=\frac{\mathcal{L}\left[S_{ee}(\mathbf{q},E)\circledast R(E)\right]}{\mathcal{L}\left[R(E)\right]}\ ,$ (3)
which is remarkably stable with respect to experimental noise Dornheim et al. (2023b, 2024a). Eq. (3) thus gives one direct access to $F_{ee}(\mathbf{q},\tau)$ from the experimental observation. As a first application, Dornheim et al. Dornheim et al. (2022) have suggested to consider the imaginary-time version of the detailed balance relation $F_{ee}(\mathbf{q},\tau)=F_{ee}(\mathbf{q},\beta-\tau)$, which gives one model-free access to the temperature of arbitrarily complex systems Dornheim et al. (2022, 2023b); Böhme et al. (2023); Schörner et al. (2023). A second practical application of the ITCF in the context of XRTS diagnostics is given by the f-sum rule, which relates the first $\tau$-derivative of $F_{ee}(\mathbf{q},\tau)$ to the momentum transfer $q=|\mathbf{q}|$ that follows from the scattering angle $\theta$ Dornheim et al. (2024a, 2023e). In this way, one can infer the normalization of the measured intensity, which is given by the electronic static structure factor $S_{ee}(\mathbf{q})$—the Fourier transform of the electron–electron pair correlation function $g_{ee}(\mathbf{r})$. Finally, we mention the recent idea by Vorberger et al. Vorberger et al. (2024), who have proposed to utilize Eq. (3) to quantify the degree of electronic non-equilibrium in the probed system.
In the present work, we extend these efforts towards a model-free diagnostics of XRTS measurements by extracting the Rayleigh weight $W_{R}(\mathbf{q})=S_{eI}^{2}(\mathbf{q})/S_{II}(\mathbf{q})$ [where $S_{eI}(\mathbf{q})$ and $S_{II}(\mathbf{q})$ are the electron–ion and ion–ion static structure factor], which describes the degree of electronic localization around the ions Vorberger and Gericke (2015); Ma et al. (2013), directly from the experimental observation. Our idea is based on a combination of the f-sum rule with the ratio of elastic to inelastic scattering $r(\mathbf{q})$, cf. Eq. (5) below. Therefore, it is generally available and even extends to non-equilibrium situations. As a practical example, we consider an XRTS measurement on strongly compressed Be that has been carried out at the NIF Döppner et al. (2023). The proposed direct inference of $W_{R}(\mathbf{q})$ helps to further constrain forward models for the interpretation of XRTS measurements, which is very important in its own right. In addition, we expect $W_{R}(\mathbf{q})$ to be a valuable diagnostic tool. For example, one might first infer the temperature $T$ from a given XRTS signal using the model-free ITCF method Dornheim et al. (2022) and then carry out a set of ab intio calculations for $W_{R}(\mathbf{q})$ over a reasonable interval of densities $\rho$; here, we use highly accurate path integral Monte Carlo (PIMC) simulations using the set-up described in Ref. Dornheim et al. (2024b) and density functional theory molecular dynamics (DFT-MD) simulations by Bethkenhagen et al. Bethkenhagen et al. . We find excellent agreement between both methods, whereas computationally less involved average-atom models Sterne et al. (2007) and Chihara fits deviate. This is of direct practical consequence for the interpretation of the experimental signal and suggests a significantly lower density of $\rho=(22\pm 2)$ g/cm3 (see also Ref. Dornheim et al. (2024b)) compared to the Chihara based estimate of $\rho=(34\pm 4)$ g/cm3 by Döppner et al. Döppner et al. (2023).
From a methodological perspective, we note that $W_{R}(\mathbf{q})$ constitutes a perfect observable for DFT-MD, as it only involves the static single-electron density distribution $n_{e}(\mathbf{r})$ [and the ion density distribution and correlation function, which are both easily accessible within MD]. This is in contrast to time-dependent DFT (TD-DFT) calculations that, in practice, contain additional approximations such as the unknown dynamic exchange–correlation kernel in the case of linear-response TD-DFT Ullrich (2012); Moldabekov et al. (2023a, b); Schörner et al. (2023). Finally, we mention the possibility to use experimental results for $W_{R}(\mathbf{q})$ as a rigorous benchmark for simulations and theoretical models in situations where the density and temperature are already being known by other means.
Idea. In general, we can split the DSF into a quasi-elastic contribution $S_{\textnormal{el}}(\mathbf{q},E)$ that is due to the electronic localization around the ions described by the Rayleigh weight, and the inelastic part $S_{\textnormal{inel}}(\mathbf{q},E)$ [cf. Fig. 1a)],
$S_{ee}(\mathbf{q},E)$ $=$ $\underbrace{W_{R}(\mathbf{q})\delta(E)}_{S_{\textnormal{el}}(\mathbf{q},E)}+S_{\textnormal{inel}}(\mathbf{q},E)\ .$ (4)
Here, the quasi-elastic nature of the first part is due to the substantially longer ionic time-scales. In practice, treating $S_{\textnormal{el}}(\mathbf{q},E)$ as a delta distribution is appropriate if the SIF $R(E)$ is significantly broader than the actual ion feature. This is usually the case except for dedicated experiments that aim to resolve ionic energy scales on purpose Descamps et al. (2020). Within the chemical picture that assumes a decomposition into effectively bound and free electrons Glenzer and Redmer (2009); Gregori et al. (2003); Kraus et al. (2019); Döppner et al. (2023), the inelastic part $S_{\textnormal{inel}}(\mathbf{q},E)$ consists of both bound-free transitions (and their reverse process Böhme et al. (2023)) and free-free transitions. The Rayleigh weight thus constitutes an indispensable ingredient to Chihara models Chihara (1987); Gregori et al. (2003) that are widely used for the interpretation of XRTS experiments with WDM Böhme et al. (2023); Döppner et al. (2023); Kraus et al. (2019); Glenzer and Redmer (2009). However, it is important to note that we do not have to make such an approximate distinction between bound and free electrons in the present work.
Let us next consider the ratio of the elastic and inelastic contributions
$r(\mathbf{q})=\frac{\int_{-\infty}^{\infty}\textnormal{d}\omega\ S_{\textnormal{el}}(\mathbf{q},\omega)}{\int_{-\infty}^{\infty}\textnormal{d}\omega\ S_{\textnormal{inel}}(\mathbf{q},\omega)}=\frac{W_{R}(\mathbf{q})}{S_{ee}(\mathbf{q})-W_{R}(\mathbf{q})}\ ,$ (5)
that constitutes a standard observable in XRTS experiments Döppner et al. (2023), see, e.g., Fig. 1a). Here $S_{ee}(\mathbf{q})=F_{ee}(\mathbf{q},0)$ denotes the aforementioned electronic static structure factor that we can also directly extract from the XRTS signal using the f-sum rule applied in the imaginary-time domain as it has been explained in detail in the recent Ref. Dornheim et al. (2024a), see also Eq. (7) below. Solving Eq. (5) for the Rayleigh weight then gives
$W_{R}(\mathbf{q})=\frac{S_{ee}(\mathbf{q})}{1+r^{-1}(\mathbf{q})}\ .$ (6)
Figure 2: Comparison of simulation results for the Rayleigh weight at $\rho=20\,$g/cc [top] and $\rho=30\,$g/cc [bottom]. Blue cross: experiment Döppner et al. (2023); solid green: PIMC; dashed red: average atom model; black and yellow dotted: chemical model with ionization states of $Z=3.52$ and $Z=3.73$; dash-dotted purple: DFT-MD Bethkenhagen et al. . PIMC and CM have been computed for $T=155.5\,$eV, DFT-MD and AA for $T=150\,$eV.
Figure 3: Curves: DFT-MD results for $W_{R}(\mathbf{q})$ at $T=150\,$eV for different mass densities $\rho$. The experimental measurement (blue cross) at the NIF Döppner et al. (2023) is associated with a density of $\rho=22\pm 2\,$g/cc [dash-dotted pink], see also the inset showing a magnified segment.
Results. In Fig. 1, we apply our new idea to an XRTS data set that has been recently obtained at the NIF by Döppner et al. Döppner et al. (2023). Panel a) shows the full XRTS intensity [solid green, cf. Eq. (1)], which can easily be decomposed into its elastic (grey) and inelastic (red) contribution if the SIF $R(E)$ is known; this is indeed the case for the backlighter x-ray source utilized at the NIF MacDonald et al. (2021). The second ingredient to Eq. (6) is given by the electronic static structure factor $S_{ee}(\mathbf{q})$, which we obtain from the Laplace transform of the XRTS signal via the f-sum rule. This procedure is illustrated in Fig. 1b), where we show the deconvolved $F_{ee}(\mathbf{q},\tau)$ as a function of $\tau$. In the imaginary-time domain, the f-sum rule is given by Dornheim et al. (2023e, c, 2024a)
$\left.\frac{\partial}{\partial\tau}F_{ee}(\mathbf{q},\tau)\right|_{\tau=0}=-\frac{\hbar^{2}\mathbf{q}^{2}}{2m_{e}}\ ;$ (7)
matching Eq. (7) [dashed red line] with the LHS. of Eq. (3) [solid blue line] around $\tau=0$ then determines the a-priori unknown normalization constant and, in turn, $S_{ee}(\mathbf{q})=F_{ee}(\mathbf{q},0)$. Inserting $r(\mathbf{q})$ and $S_{ee}(\mathbf{q})$ into Eq. (6) gives our final, model-free estimate of the Rayleigh weight, which is shown as the blue cross in Fig. 1c) at the corresponding experimental wavenumber of $q=7.89\,$Å-1. The red, green and yellow solid curves in the panel show highly accurate ab initio PIMC results that have been obtained for $N_{\textnormal{Be}}=10$ Be atoms using the set-up described in the recent Refs. Dornheim et al. (2024b, c) at the temperature of $T=155.5$ eV that has been inferred from the symmetry of the ITCF; see the appendix of Ref. Dornheim et al. (2024b) for additional details. In addition, the associated colored areas correspond to PIMC simulations for $T=(155.5\pm 15)$ eV and indicate the uncertainty in the Rayleigh weight at a given density due to the uncertainty in the inferred temperature. Fig. 1 thus indicates a mass density of $\rho\approx 20$ g/cm3, whereas the nominal result of $\rho=(34\pm 4)$ g/cm3 reported in the original Ref. Döppner et al. (2023) is ruled out.
To get further insights into the sensitivity of the interpretation of a given XRTS measurement to the employed model or simulation technique, we compare a number of independent methods in Fig. 2, where the top and bottom panels correspond to $\rho=20$ g/cm3 and $\rho=30$ g/cm3, respectively. For both cases, we find very good agreement between PIMC (solid green) and DFT-MD simulations by Bethkenhagen et al. Bethkenhagen et al. (dash-dotted purple). The computationally cheaper average-atom model Sterne et al. (2007) (dashed red) is in qualitative, though not quantitative agreement with the ab initio data sets. Instead, it agrees with the experimental data point for $\rho=30$ g/cm3. This clearly highlights the sensitivity of equation-of-state measurements to the proper treatment of electronic correlations and other many-body effects in the analysis of the experimental observation. Finally, we include two chemical models for the hypothetical ionization degrees of $Z=3.52$ (dotted black) and $Z=3.73$ (dotted yellow) for both densities, see the Supplemental Material for additional details sup . Evidently, the inferred density strongly depends on the inferred ionization state, making the invocation of additional constraints desirable. In Ref. Döppner et al. (2023), Döppner et al. have used the Saha equation in combination with a semi-empirical form factor lowering model for this purpose, leading to the nominal parameters of $\rho=(34\pm 4)$ g/cm3 and $Z=3.4\pm 0.1$. A systematic investigation of the origin of this discrepancy between their chemical model and the ab initio data sets from PIMC and DFT-MD is beyond the scope of the present work and will be pursued in a dedicated future study.
Let us conclude with a systematic analysis of the dependence of $W_{R}(\mathbf{q})$ on the density that is shown in Fig. 3 based on extensive DFT-MD results for $\rho=(5-80)$ g/cm3 at $T=150$ eV. The main trend is given by the systematic increase of the Rayleigh weight with $\rho$, which is to a large degree due to the different length scales in the system; the effect almost vanishes when one adjusts the $x$-axis by the Fermi wavenumber $q_{\textnormal{F}}\sim 1/\rho^{1/3}$. The inset shows a magnified segment around the experimental data point, with the latter being located between the DFT-MD results for $\rho=20$ g/cm3 (red) and $\rho=24$ g/cm3 (black); this leads to our final estimate for the density of $\rho=(22\pm 2)$ g/cm3 [dash-dotted pink].
Summary and Discussion. We have presented a new approach for the model-free extraction of the Rayleigh weight $W_{R}(\mathbf{q})$ from XRTS measurements. Most importantly, $W_{R}(\mathbf{q})$ constitutes an important measure for the electronic localization around the ions, which is interesting in its own right. In addition, $W_{R}(\mathbf{q})$ is of direct practical value for XRTS diagnostics, as we have demonstrated in detail on a recent experiment with strongly compressed beryllium at the NIF Döppner et al. (2023). In particular, we propose to first infer the temperature $T$ from the model-free ITCF thermometry approach Dornheim et al. (2022, 2023b), and to subsequently match the experimental result for $W_{R}(\mathbf{q})$ with simulation results over a reasonable range of densities. We note that $W_{R}(\mathbf{q})$ is a particularly suitable observable for DFT-MD simulations, as it does not involve any dynamic information such as the a-priori unknown dynamic exchange–correlation kernel. We find excellent agreement between DFT-MD and PIMC reference data; this is encouraging since direct PIMC simulations are still limited to low-$Z$ materials at moderate to high temperatures Dornheim (2019); Bonitz et al. (2024), whereas DFT-MD is more broadly applicable. An additional advantage of $W_{R}(\mathbf{q})$ over alternative observables such as the ratio of elastic to inelastic scattering $r(\mathbf{q})$ is that the former does not require any explicit information about the electronic static structure factor $S_{ee}(\mathbf{q})$, which is notoriously difficult for DFT-based methodologies.
In practice, we infer a mass density of $\rho=(22\pm 2)$ g/cm3 from the beryllium data set using either PIMC or DFT-MD simulations, which is significantly lower than the nominal value of $\rho=(34\pm 4)$ g/cm3 that has been reported in the original publication based on a chemical Chihara model. Moreover, we find that the computationally cheaper average-atom model tends to overestimate the density, whereas chemical models generally require additional constraints for parameters such as the ionization degree. Our study thus clearly highlights the importance of an accurate treatment of quantum many-body effects for XRTS based equation-of-state measurements even at relatively high temperatures; the main value of average-atom calculations is their low computational efforts, making them a useful tool to constrain the parameter space for more expensive ab initio simulations.
We are convinced that our work opens up new possibilities for the study of warm dense matter and beyond. An important point for future research is given by the sensitivity of the inferred Rayleigh weight to the SIF $R(E)$, which is usually modelled for backlighter set-ups MacDonald et al. (2021), but can be known with high precision at modern XFEL facilities Gawne et al. (2024a). Indeed, recent advances in high-resolution XRTS measurements Gawne et al. (2024b) at the European XFEL will likely facilitate the application of the model-free ITCF thermometry technique even at moderate temperatures of $T\sim 1\,$eV, which are of relevance for both planetary and material science; the current model-free framework for the extraction of $W_{R}(\mathbf{q})$ is applicable at any temperature. In combination, these two methods will allow for accurate equation-of-state measurements in previously inaccessible regimes. Finally, we mention the intriguing possibility of performing an XRTS experiment on an isochorically heated sample with an appropriate delay between pump and probe to ensure proper equilibration. Since both $T$ and $\rho$ would be well known in such a scenario, a corresponding measurement of $W_{R}(\mathbf{q})$ using the present framework would constitute a truly unambiguous reference data set for the rigorous benchmarking of ab initio simulations and chemical models alike.
Acknowledgments
This work was partially supported by the Center for Advanced Systems Understanding (CASUS), financed by Germany’s Federal Ministry of Education and Research (BMBF) and the Saxon state government out of the State budget approved by the Saxon State Parliament. This work has received funding from the European Union’s Just Transition Fund (JTF) within the project Röntgenlaser-Optimierung der Laserfusion (ROLF), contract number 5086999001, co-financed by the Saxon state government out of the State budget approved by the Saxon State Parliament. This work has received funding from the European Research Council (ERC) under the European Union’s Horizon 2022 research and innovation programme (Grant agreement No. 101076233, ”PREXTREME”). This work was partially supported by the German Research Foundation (DFG) within the Research Unit FOR 2440. Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. Computations were performed on a Bull Cluster at the Center for Information Services and High-Performance Computing (ZIH) at Technische Universität Dresden and at the Norddeutscher Verbund für Hoch- und Höchstleistungsrechnen (HLRN) under grants mvp00018 and mvp00024. SH was supported by Sandia National Laboratories, a multi-mission laboratory managed and operated by National Technology and Engineering Solutions of Sandia, LLC, a wholly owned subsidiary of Honeywell International, Inc., for DOE’s National Nuclear Security Administration under contract DE-NA0003525. This paper describes objective technical results and analysis. Any subjective views or opinions that might be expressed in the paper do not necessarily represent the views of the U.S. Department of Energy or the United States Government.

### v2

@@
 In principle, it gives one access to important system parameters such as the temperature, density, and ionization state, but the interpretation of the measured XRTS intensity usually relies on theoretical models and approximations.
-In this work, we show that it is possible to extract the Rayleigh weight—a key property that describes the electronic localization around the ions—directly from the experimental data without the need for any model calculations or simulations.
+In this context, a key property is given by the Rayleigh weight that describes the electronic localization around the ions.
+Here, we show that it is possible to extract the Rayleigh weight directly from the experimental data without the need for any model calculations or simulations.
 As a practical application, we consider an experimental measurement of strongly compressed Be at the National Ignition Facility (NIF) [Döppner et al., Nature 618, 270-275 (2023)].
-In addition to being interesting in their own right, our results will open up new avenues for diagnostics from ab initio simulations, help to further constrain existing chemical models, and constitute a rigorous benchmark for theory and simulations.
+We demonstrate that experimental results for the Rayleigh weight open up new avenues for the interpretation of XRTS experiments by matching the measurement with ab initio simulations such as density functional theory or path integral Monte Carlo.
+Interestingly, this new procedure leads to significantly lower density compared to previously used chemical models.
 The x-ray Thomson scattering (XRTS) method Glenzer and Redmer (2009); Sheffield et al.
@@
 (1983).
-Moreover, WDM plays an important role in a variety of technological applications such as inertial confinement fusion (ICF) Betti and Hurricane (2016), where the fuel capsule has to traverse the WDM regime in a controlled way to reach ignition Hu et al.
+Moreover, WDM plays an important role in a variety of technological applications such as inertial confinement fusion (ICF) Betti and Hurricane (2016), where the fuel capsule has to traverse the WDM regime in a controlled way to reach ignition ($\rho=500$ g/cc, $T=5\times 10^{7}$ K) Hu et al.
 (2011).
 The recent breakthrough at the National Ignition Facility (NIF) to reach ignition The Indirect Drive ICF Collaboration (2024) has further substantiated the importance of accurately diagnosing WDM states.
-In the laboratory, these extreme states can be realized using a variety of experimental techniques Falk (2018), and XRTS is often used to infer a-priori unknown system parameters such as the mass density $\rho$, temperature $T$, and ionization state $Z$; see, e.g., Ref.
+In the laboratory, these extreme states can be created in the laboratory using a variety of experimental techniques Falk (2018), and XRTS is often used to infer a-priori unknown system parameters such as the mass density $\rho$, temperature $T$, and ionization state $Z$; see, e.g., Ref.
 Plagemann et al.
@@
 Naturally, the thus inferred information depends on the employed forward model for $S_{ee}(\mathbf{q},E)$.
-These are usually based on a number of assumptions such as the possibility to distinguish between bound and free electrons as proposed within the popular Chihara approach Chihara (1987); Höll et al.
+Although advances with ab-initio methods have been made Baczewski et al.
+(2016), these models are usually based on a number of assumptions such as the possibility to distinguish between bound and free electrons as proposed within the popular Chihara approach Chihara (1987); Höll et al.
+(2007); Glenzer et al.
+(2007); Lee et al.
+(2009, 2010); Döppner et al.
 (2010); Fortmann et al.
-(2009); Glenzer and Redmer (2009); Kraus et al.
+(2009); Glenzer and Redmer (2009); Gregori et al.
+(2003); Johnson et al.
+(2012); Souza et al.
+(2014); Mattern et al.
+(2012); Kraus et al.
 (2019); Döppner et al.
@@
 (2023c, d).
-It corresponds to the usual intermediate scattering function $F(\mathbf{q},t)=\braket{\hat{n}(\mathbf{q},t)\hat{n}(-\mathbf{q},0)}$ with an imaginary argument $t=-i\hbar\tau$ with $\tau\in[0,\beta]$ and $\beta=1/k_{\textnormal{B}}T$ the inverse temperature.
+It corresponds to the usual intermediate scattering function $F(\mathbf{q},t)=\braket{\hat{n}(\mathbf{q},t)\hat{n}(-\mathbf{q},0)}$ with an imaginary time argument $t=-i\hbar\tau$ with $\tau\in[0,\beta]$ and $\beta=1/k_{\textnormal{B}}T$ the inverse temperature.
 A key advantage of working in the Laplace domain is given by the convolution theorem
@@
 A second practical application of the ITCF in the context of XRTS diagnostics is given by the f-sum rule, which relates the first $\tau$-derivative of $F_{ee}(\mathbf{q},\tau)$ to the momentum transfer $q=|\mathbf{q}|$ that follows from the scattering angle $\theta$ Dornheim et al.
-(2024a, 2023e).
+(2024a, 2023e), see Eq.
+(7), with $m_{e}$ the electron mass.
 In this way, one can infer the normalization of the measured intensity, which is given by the electronic static structure factor $S_{ee}(\mathbf{q})$—the Fourier transform of the electron–electron pair correlation function $g_{ee}(\mathbf{r})$.
@@
 .
-We find excellent agreement between both methods, whereas computationally less involved average-atom models Sterne et al.
-(2007) and Chihara fits deviate.
-This is of direct practical consequence for the interpretation of the experimental signal and suggests a significantly lower density of $\rho=(22\pm 2)$ g/cm3 (see also Ref.
+We find excellent agreement between both methods in $W_{R}(\mathbf{q})$, whereas computationally less intensive average-atom models Sterne et al.
+(2007); Souza et al.
+(2014) and parameterized Chihara fits Gregori et al.
+(2003); Glenzer and Redmer (2009) deviate.
+This is of direct practical consequence for the interpretation of the experimental signal and suggests a significantly lower density of $\rho=(22\pm 2)$ g/cc (see also Ref.
 Dornheim et al.
-(2024b)) compared to the Chihara based estimate of $\rho=(34\pm 4)$ g/cm3 by Döppner et al.
+(2024b)) compared to the Chihara based estimate of $\rho=(34\pm 4)$ g/cc by Döppner et al.
 Döppner et al.
@@
 From a methodological perspective, we note that $W_{R}(\mathbf{q})$ constitutes a perfect observable for DFT-MD, as it only involves the static single-electron density distribution $n_{e}(\mathbf{r})$ [and the ion density distribution and correlation function, which are both easily accessible within MD].
-This is in contrast to time-dependent DFT (TD-DFT) calculations that, in practice, contain additional approximations such as the unknown dynamic exchange–correlation kernel in the case of linear-response TD-DFT Ullrich (2012); Moldabekov et al.
+This is in contrast to time-dependent DFT (TD-DFT) calculations that, in practice, contain additional approximations such as the unknown dynamic exchange–correlation kernel in the case of linear-response TD-DFT Baczewski et al.
+(2016); Moldabekov et al.
 (2023a, b); Schörner et al.
 (2023).
-Finally, we mention the possibility to use experimental results for $W_{R}(\mathbf{q})$ as a rigorous benchmark for simulations and theoretical models in situations where the density and temperature are already being known by other means.
+Finally, we mention the possibility to use experimental results for $W_{R}(\mathbf{q})$ as a rigorous benchmark for simulations and theoretical models in situations where the density and temperature are already known by other means.
 Idea.
-In general, we can split the DSF into a quasi-elastic contribution $S_{\textnormal{el}}(\mathbf{q},E)$ that is due to the electronic localization around the ions described by the Rayleigh weight, and the inelastic part $S_{\textnormal{inel}}(\mathbf{q},E)$ [cf.
+In general, when we can distinguish an elastic peak in an XRTS signal, we can split the DSF into a quasi-elastic contribution $S_{\textnormal{el}}(\mathbf{q},E)$ described by the Rayleigh weight, and an inelastic part $S_{\textnormal{inel}}(\mathbf{q},E)$ [cf.
 Fig.
@@
 $S_{ee}(\mathbf{q},E)$ $=$ $\underbrace{W_{R}(\mathbf{q})\delta(E)}_{S_{\textnormal{el}}(\mathbf{q},E)}+S_{\textnormal{inel}}(\mathbf{q},E)\ .$ (4)
-Here, the quasi-elastic nature of the first part is due to the substantially longer ionic time-scales.
+Here, the quasi-elastic nature of the first part is due to the substantially longer ionic time-scales and localization of bound and nearly-bound electrons around the ions.
 In practice, treating $S_{\textnormal{el}}(\mathbf{q},E)$ as a delta distribution is appropriate if the SIF $R(E)$ is significantly broader than the actual ion feature.
-This is usually the case except for dedicated experiments that aim to resolve ionic energy scales on purpose Descamps et al.
+This is usually the case except for dedicated experiments that aim to resolve ionic energy scales  Descamps et al.
+(2020); White et al.
 (2024).
@@
 (2019); Döppner et al.
-(2023), the inelastic part $S_{\textnormal{inel}}(\mathbf{q},E)$ consists of both bound-free transitions (and their reverse process Böhme et al.
-(2023)) and free-free transitions.
-The Rayleigh weight thus constitutes an indispensable ingredient to Chihara models Chihara (1987); Gregori et al.
+(2023), the inelastic part $S_{\textnormal{inel}}(\mathbf{q},E)$ consists of both transitions between bound and free states (and their reverse process Böhme et al.
+(2023)) and transitions between two free electronic states.
+The Rayleigh peak, too, is informed by both free electrons (screening cloud) and bound electrons (form factor), and constitutes an indispensable ingredient to Chihara models Chihara (1987); Gregori et al.
+(2003); Souza et al.
 (2014) that are widely used for the interpretation of XRTS experiments with WDM Böhme et al.
@@
 (2019); Glenzer and Redmer (2009).
-However, it is important to note that we do not have to make such an approximate distinction between bound and free electrons in the present work.
+However, it is important to note that we do not have to make such an approximate distinction between bound and free electrons in the present work as the computation of $W_{R}(q)$ only involves correlation functions between all electrons and ions, i.e., $S_{eI}(q)$ and $S_{II}(q)$.
 Let us next consider the ratio of the elastic and inelastic contributions
@@
 1a).
-Here $S_{ee}(\mathbf{q})=F_{ee}(\mathbf{q},0)$ denotes the aforementioned electronic static structure factor that we can also directly extract from the XRTS signal using the f-sum rule applied in the imaginary-time domain as it has been explained in detail in the recent Ref.
+Here $S_{ee}(\mathbf{q})=F_{ee}(\mathbf{q},0)$ denotes the aforementioned electronic static structure factor that we can also directly extract from the XRTS signal using the f-sum rule applied in the imaginary-time domain as has been explained in detail in the recent Ref.
 Dornheim et al.
@@
 1c) at the corresponding experimental wavenumber of $q=7.89\,$Å-1.
-The red, green and yellow solid curves in the panel show highly accurate ab initio PIMC results that have been obtained for $N_{\textnormal{Be}}=10$ Be atoms using the set-up described in the recent Refs.
+The red, green and yellow solid curves in the panel show ab initio PIMC results that have been obtained for $N_{\textnormal{Be}}=10$ Be atoms using the set-up described in the recent Refs.
 Dornheim et al.
@@
 Fig.
-1 thus indicates a mass density of $\rho\approx 20$ g/cm3, whereas the nominal result of $\rho=(34\pm 4)$ g/cm3 reported in the original Ref.
+1 thus indicates a mass density of $\rho\approx 20$ g/cc, whereas the nominal result of $\rho=(34\pm 4)$ g/cc reported in the original Ref.
 Döppner et al.
 (2023) is ruled out.
-To get further insights into the sensitivity of the interpretation of a given XRTS measurement to the employed model or simulation technique, we compare a number of independent methods in Fig.
-2, where the top and bottom panels correspond to $\rho=20$ g/cm3 and $\rho=30$ g/cm3, respectively.
+For further insight into the sensitivity of the interpretation of a given XRTS measurement to the employed model or simulation technique, we compare a number of independent methods in Fig.
+2, where the top and bottom panels correspond to $\rho=20$ g/cc and $\rho=30$ g/cc, respectively.
 For both cases, we find very good agreement between PIMC (solid green) and DFT-MD simulations by Bethkenhagen et al.
@@
 The computationally cheaper average-atom model Sterne et al.
+(2007); Souza et al.
 (2014) (dashed red) is in qualitative, though not quantitative agreement with the ab initio data sets.
-Instead, it agrees with the experimental data point for $\rho=30$ g/cm3.
-This clearly highlights the sensitivity of equation-of-state measurements to the proper treatment of electronic correlations and other many-body effects in the analysis of the experimental observation.
+Instead, it agrees with the experimental data point for $\rho=30$ g/cc.
+This highlights the sensitivity of equation-of-state measurements to the proper treatment of electronic correlations and other many-body effects in the analysis of the experimental observation.
 Finally, we include two chemical models for the hypothetical ionization degrees of $Z=3.52$ (dotted black) and $Z=3.73$ (dotted yellow) for both densities, see the Supplemental Material for additional details sup .
@@
 (2023), Döppner et al.
-have used the Saha equation in combination with a semi-empirical form factor lowering model for this purpose, leading to the nominal parameters of $\rho=(34\pm 4)$ g/cm3 and $Z=3.4\pm 0.1$.
+have used the Saha equation in combination with a semi-empirical form factor lowering model for this purpose, leading to the nominal parameters of $\rho=(34\pm 4)$ g/cc and $Z=3.4\pm 0.1$.
 A systematic investigation of the origin of this discrepancy between their chemical model and the ab initio data sets from PIMC and DFT-MD is beyond the scope of the present work and will be pursued in a dedicated future study.
-Let us conclude with a systematic analysis of the dependence of $W_{R}(\mathbf{q})$ on the density that is shown in Fig.
-3 based on extensive DFT-MD results for $\rho=(5-80)$ g/cm3 at $T=150$ eV.
+Let us conclude with a systematic analysis of the dependence of $W_{R}(\mathbf{q})$ on the density, as shown in Fig.
+3 based on extensive DFT-MD results for $\rho=(5-80)$ g/cc at $T=150$ eV.
 The main trend is given by the systematic increase of the Rayleigh weight with $\rho$, which is to a large degree due to the different length scales in the system; the effect almost vanishes when one adjusts the $x$-axis by the Fermi wavenumber $q_{\textnormal{F}}\sim 1/\rho^{1/3}$.
-The inset shows a magnified segment around the experimental data point, with the latter being located between the DFT-MD results for $\rho=20$ g/cm3 (red) and $\rho=24$ g/cm3 (black); this leads to our final estimate for the density of $\rho=(22\pm 2)$ g/cm3 [dash-dotted pink].
+The inset shows a magnified segment around the experimental data point, with the latter being located between the DFT-MD results for $\rho=20$ g/cc (red) and $\rho=24$ g/cc (black); this leads to our final estimate for the density of $\rho=(22\pm 2)$ g/cc [dash-dotted pink].
 Summary and Discussion.
 We have presented a new approach for the model-free extraction of the Rayleigh weight $W_{R}(\mathbf{q})$ from XRTS measurements.
-Most importantly, $W_{R}(\mathbf{q})$ constitutes an important measure for the electronic localization around the ions, which is interesting in its own right.
+Most importantly, $W_{R}(\mathbf{q})$ constitutes an important measure for the electronic localization around the ions, which is related to multiple possible definitions of the average ion charge Murillo et al.
+(2013) and interesting in its own right.
 In addition, $W_{R}(\mathbf{q})$ is of direct practical value for XRTS diagnostics, as we have demonstrated in detail on a recent experiment with strongly compressed beryllium at the NIF Döppner et al.
@@
 We find excellent agreement between DFT-MD and PIMC reference data; this is encouraging since direct PIMC simulations are still limited to low-$Z$ materials at moderate to high temperatures Dornheim (2019); Bonitz et al.
-(2024), whereas DFT-MD is more broadly applicable.
+(2024), whereas DFT-MD and TD-DFT Baczewski et al.
+(2016); Gawne et al.
+(2024b) are more broadly applicable.
 An additional advantage of $W_{R}(\mathbf{q})$ over alternative observables such as the ratio of elastic to inelastic scattering $r(\mathbf{q})$ is that the former does not require any explicit information about the electronic static structure factor $S_{ee}(\mathbf{q})$, which is notoriously difficult for DFT-based methodologies.
-In practice, we infer a mass density of $\rho=(22\pm 2)$ g/cm3 from the beryllium data set using either PIMC or DFT-MD simulations, which is significantly lower than the nominal value of $\rho=(34\pm 4)$ g/cm3 that has been reported in the original publication based on a chemical Chihara model.
-Moreover, we find that the computationally cheaper average-atom model tends to overestimate the density, whereas chemical models generally require additional constraints for parameters such as the ionization degree.
-Our study thus clearly highlights the importance of an accurate treatment of quantum many-body effects for XRTS based equation-of-state measurements even at relatively high temperatures; the main value of average-atom calculations is their low computational efforts, making them a useful tool to constrain the parameter space for more expensive ab initio simulations.
-We are convinced that our work opens up new possibilities for the study of warm dense matter and beyond.
+In practice, we infer a mass density of $\rho=(22\pm 2)$ g/cc from the beryllium data set using either PIMC or DFT-MD simulations, which is significantly lower than the nominal value of $\rho=(34\pm 4)$ g/cc that has been reported in the original publication based on a chemical Chihara model; note that the latter must take into account a proper multi-component treatment of differently charged ions Wünsch et al.
+(2011).
+Moreover, we find that the computationally more efficient average-atom model tends to overestimate the density, whereas chemical models generally require additional constraints for parameters such as the ionization degree.
+Our study thus clearly highlights the importance of an accurate treatment of quantum many-body effects for XRTS based equation-of-state measurements even at relatively high temperatures.
+This work opens up new possibilities for the study of warm dense matter and beyond.
 An important point for future research is given by the sensitivity of the inferred Rayleigh weight to the SIF $R(E)$, which is usually modelled for backlighter set-ups MacDonald et al.
@@
 Department of Energy or the United States Government.
+Supplemental Material
+Appendix A Uncertainty estimate of NIF data point
+As specified in Eq.
+(6) of the main text, the model-free extraction of the Rayleigh weight from the measured XRTS intensity is based on two ingredients, each of which contributes to the total uncertainty of the final result for $W_{R}(\mathbf{q})$: i) the electron–electron static structure factor $S_{ee}(\mathbf{q})$ and ii) the ratio of elastic to inelastic scattering $r(\mathbf{q})$.
+i) In principle, evaluating the two-sided Laplace transform would require an infinite integration range, whereas the spectral window in the experiment is always finite.
+In practice, we thus define the symmetrically truncated Laplace transform
+$\mathcal{L}\left[S_{ee}(\mathbf{q},E)\right]=\int_{-x}^{x}\textnormal{d}E\ S_{ee}(\mathbf{q},E)\ e^{-\tau E}\ ,$ (8)
+and investigate the convergence with respect to $x$, see Ref.
+Dornheim et al.
+(2024a) for a detailed description of its implications for the determination of $S_{ee}(\mathbf{q})$ from the f-sum rule, i.e., from the first derivative of $F_{ee}(\mathbf{q},\tau)$ with respect to $\tau$ around $\tau=0$.
+For the XRTS data set of strongly compressed beryllium that is considered in the present work, a corresponding convergence analysis is shown in Ref.
+Dornheim et al.
+(2024b), leading to $S_{ee}(\mathbf{q})=0.971\pm 0.036$.
+ii) To decompose the experimental intensity signal $I(\mathbf{q},E)$ into an elastic and inelastic contribution as it has been depicted in Fig.
+1a) of the main text, we only need to assume that $S_{\textnormal{inel}}(\mathbf{q},E)$ is smooth around the beam energy $E_{0}=9000\,$eV.
+In practice, it is convenient to use a Chihara fit onto the observed data for this purpose, which leads to $r(\mathbf{q})=0.087\pm 0.004$ Döppner et al.
+(2023); the corresponding error is obtained from a Gaussian error propagation onto the elastic-to-inelastic ratio from the given fit uncertainties.
+It is important to note that $r(\mathbf{q})$ is to a very high degree independent from a particular forward model for $S_{ee}(\mathbf{q},E)$ as long as the convolution with the SIF fits the experimental signal $I(\mathbf{q},E)$.
+This is in stark contrast to parameters such as the density, temperature, or charge state.
+We also note that while the core 1s electrons are tightly bound and well localized in the present case, other materials at other conditions may have loosely bound valence electrons that increase uncertainties in isolating the elastic contribution.
+The final uncertainty of our estimate for $W_{R}(\mathbf{q})$ then follows from a straightforward error propagation analysis of Eq. (6), and we find $W_{R}(\mathbf{q})=0.0778\pm 0.0044$.
+Appendix B Range of applicability and limitations of the method
+Being exclusively based on the ratio of elastic-to-inelastic scattering and the static structure factor, the method is applicable to a wide range of conditions.
+In particular, one only needs to resolve the ITCF $F_{ee}(\mathbf{q},\tau)$ around the vicinity of $\tau=0$.
+This is in stark contrast to the model-free ITCF thermometry approach discussed in Refs.
+Dornheim et al.
+(2022, 2023b), for which one needs to resolve the ITCF for $\tau\gtrsim\beta/2$.
+As a consequence, the temperature extraction only works for sufficiently hot systems (i.e., small enough $\beta$), whereas the present approach works well for all temperatures, including ambient conditions Dornheim et al.
+(2024a).
+A more serious point of consideration is given by the convolution theorem [Eq.
+(3) of the main text].
+Most fundamentally, this relation assumes that the SIF $R(E)$ does not depend on the photon energy.
+A dedicated recent analysis by Gawne et al.
+Gawne et al.
+(2024a) has shown that this is indeed the case to a high degree for usual experimental energy ranges, and, therefore, can be safely neglected here.
+Nevertheless, this subtlety might become important if one were to measure the XRTS spectrum over several thousand eV, as it would be required for heavier elements, see below.
+A second point concerning Eq.
+(3) is given by knowledge of $R(E)$ itself, which is indispensable for both forward-modeling based methodologies Glenzer and Redmer (2009) as well as the present model-free approach.
+While $R(E)$ can be inferred at modern XFEL facilities such as the European XFEL with high confidence Gawne et al.
+(2024a), its characterization for backlighter sources like the one utilized for the XRTS measurement on strongly compressed beryllium at the NIF Döppner et al.
+(2023) is more difficult MacDonald et al.
+(2021).
+Clearly, any uncertainty related to $R(E)$ directly translates to the model-free estimate of $W_{R}(\mathbf{q})$ using the present approach, and might thus influence the extracted system parameters such as the mass density $\rho$.
+Hence, we strongly encourage dedicated measurements for the characterization of $R(E)$ for any given experimental set-up if possible.
+In addition to these technical considerations, other potential sources of uncertainty concern the probed state of matter itself.
+A well-known issue related to the experimental realization of warm dense matter states Falk (2018) is given by possible non-equilibrium effects.
+These can potentially be detected without the need for models from the ITCF itself, see the recent Ref.
+Vorberger et al.
+(2024).
+A second issue that might be important – in particular for spherical implosion experiments – is given by the potential inhomogeity of the compressed sample Chapman et al.
+(2014); Döppner et al.
+(2023), which deserves dedicated attention in future studies.
+A third concern is related to the treatment of loosely bound or “hopping” electronic states, which have some properties of bound states (localization near the ions) and some properties of free states (extensive tails).
+This concern is related to the Chihara distinction between free and bound electrons Murillo et al.
+(2013).
+Finally, we mention the importance of resolving all important spectral features in the XRTS measurement for a given material; this is important both for the determination of $r(\mathbf{q})$ and of $S_{ee}(\mathbf{q})$ (though not for the ITCF thermometry approach Dornheim et al.
+(2023b)), and requires increasingly large spectral ranges with increasing nuclear charge.
+Appendix C Path integral Monte Carlo
+As written in the main text, the Rayleigh weight is easily computed from the electron–ion and ion–ion static structure factors via Vorberger and Gericke (2015)
+$W_{R}(\mathbf{q})=\frac{S_{eI}^{2}(\mathbf{q})}{S_{II}(\mathbf{q})}\ .$ (9)
+Both static structure factors are directly accessible to PIMC simulations; we use the approach presented in the recent Refs.
+Dornheim et al.
+(2024b, c) using $P=200$ imaginary-time propagators where the divergent electron–ion Coulomb attraction is treated via the pair approximation as described in Refs.
+Böhme et al.
+(2023); Militzer (2016).
+All PIMC simulations are carried out with the ISHTAR code Dornheim et al.
+(2024d), using the canonical adaption of the worm algorithm by Boninsegni et al.
+Boninsegni et al.
+(2006a, b) that has been introduced in Ref.
+Dornheim et al.
+(2021).
+Appendix D Density functional theory molecular dynamics
+DFT-MD simulations give snapshots of the ion positions in the simulation box as well as the electronic density on a grid in the same supercell.
+These can be used to calculate the ion-ion structure factor
+$S_{II}(\mathbf{q})=\frac{1}{N_{I}}\langle\rho_{I}(\mathbf{q})\rho_{I}^{*}(\mathbf{q})\rangle\ ,$ (10)
+as well as the electron-ion static structure factor
+$S_{eI}(\mathbf{q})=\frac{1}{\sqrt{N_{e}N_{I}}}\langle\rho_{e}(\mathbf{q})\rho_{I}^{*}(\mathbf{q})\rangle\ ,$ (11)
+from which the Rayleigh weight derives Vorberger and Gericke (2015).
+The DFT-MD simulations along the 150 eV isotherm were performed with the Vienna Ab initio Simulation Package (VASP) Kresse and Hafner (1993, 1994); Kresse and Furthmüller (1996) and were partially taken from Refs.
+Döppner et al.
+(2023); Bethkenhagen et al.
+.
+Here, we considered 11 densities ranging from 5 g/cc to 80 g/cc, whereas the calculations were carried out with 888864646464 beryllium atoms in the simulation box depending on the simulated density.
+All electrons were treated explicitly using the Coulomb potential with a cutoff energy of 10 keV.
+The DFT-MD simulations were typically run for at least 20 000 time steps after equilibration with a time step size between 5 as and 50 as.
+The ion temperature was controlled by a Nosé-Hoover thermostat Nosé (1984), the Brillouin zone was sampled at the Baldereschi mean value point Baldereschi (1973), and we employed the exchange-correlation (XC) functional of Perdew, Burke, and Ernzerhof (PBE) Perdew et al.
+(1996).
+Note, that the ion density can be calculated in every timestep, while the electron density was only obtained for 15-20 snapshots per DFT-MD simulation.
+Therefore, the ion-ion structure factor converges much faster than the electron-ion static structure factor.
+Figure 4: Ion–ion static structure factor $S_{II}(q)$ [left] and electron–ion static structure factor $S_{eI}(q)$ [right] for $\rho=20\,$g/cc and $T\sim 150\,$eV.
+The green crosses and dashed red lines correspond to PIMC (taken from Ref.
+Dornheim et al.
+(2024b)) and average atom, respectively and the yellow (black) dotted curve to the chemical model with $Z=3.73$ ($Z=3.52$).
+Appendix E Average atom model
+Average-atom models Liberman (1979); Wilson et al.
+(2006); Sterne et al.
+(2007) solve for the real-space electron density around a single ion, using density functional theory to obtain all-electron, spherically averaged densities $n_{e}^{\mathrm{full}}(r)$ and self-consistent radial potentials within an ion sphere defined by the Wigner-Seitz radius.
+The model used here is a semi-relativistic version of the Purgatorio code Wilson et al.
+(2006); Sterne et al.
+(2007) that has been extended following Starrett and Saumon Starrett and Saumon (2014) to compute the response of the background plasma by solving for the external electron density $n_{e}^{\mathrm{ext}}(r)$ in the ion sphere without a central nuclear charge.
+The total form factor $f(q)$ is given by the Fourier transform of the pseudo-atom density $n_{e}^{\mathrm{PA}}(r)=n_{e}^{\mathrm{full}}(r)-n_{e}^{\mathrm{ext}}(r)$.
+Note that since $\int\,n_{e}^{\mathrm{PA}}(r)d^{3}r=Z_{n}$, with $Z_{n}$ the nuclear charge, $f(0)=Z_{n}$.
+A screening density is defined by subtracting the bound electron density from $n_{e}^{\mathrm{PA}}(r)$.
+The transform of the screening density is combined with a response function to generate an ion-ion potential that determines the ion structure factor $S_{II}(q)$.
+Finally, the Rayleigh weight is obtained via:
+$W_{R}(q)=\frac{f^{2}(q)S_{II}(q)}{Z_{n}}\ .$ (12)
+Here, we use Kohn-Sham LDA for the exchange potential and calculate the ion charge by integrating the free-electron density of states multiplied by the Fermi occupation factor with a self-consistent chemical potential.
+For Be at 20 g/cc and 150 eV, we find an average ionization of 3.1 and relatively minor ($\approx$10%) variations of $W_{R}(q)$ with alternative exchange potentials and definitions of the average ion charge.
+With this AA model, XRTS spectra can also be directly computed following Johnson et al.
+(2012); Souza et al.
+(2014).
+Finally, we note that using the Fourier transform of the average-atom electron density $n_{e}^{\mathrm{full}}(r)$ for $f(q)$ can give oscillations inconsistent with the $W_{R}(q)$ of PIMC and DFT-MD.
+Appendix F Chihara models
+The Chihara decomposition Chihara (1987) is used to separate out contributions from bound and free electrons.
+Thus, in this picture, the Rayleigh weight is calculated by considering the ionic contributions.
+Figure 5: DFT-MD results for the ion–ion static structure factor $S_{II}(q)$ at $T=150\,$eV and $\rho=5\,$g/cc for $N_{Be}=8$ (black) and $N_{Be}=20$ (red) Be atoms.
+More specifically, for a multi-component system with species $a$ and $b$, the Rayleigh weight is calculated as Wünsch et al. (2011):
+$W_{R}(q)=\sum_{a,b}\Bigl{\{}\sqrt{x_{a}x_{b}}\left[f_{\text{a}}^{\text{ff}}(q)+f_{\text{a}}^{\text{sc}}(q)\right]\\ \times\left[f_{\text{b}}^{\text{ff}}(q)+f_{\text{b}}^{\text{sc}}(q)\right]S_{ab}(q)\Bigr{\}}\,.$ (13)
+with $q=|\mathbf{q}|$ and $x_{a}$ being the partial number density fraction of ions of species $a$ given by
+$x_{a}=n_{a}/n_{i}$ (14)
+for the mean number density $n_{a}$ of species $a$. The total number density of ions (with the mean ion mass) in the plasma is
+$n_{i}=\sum_{a}n_{a}\,.$ (15)
+Here, the ionic form factor for a given species, e.g., $a$, is defined as $f_{\text{a}}^{\text{ff}}(q)$ and $f_{\text{a}}^{\text{sc}}(q)$ denotes the respective screening cloud.
+Partial static structure factors $S_{ab}$ account for correlations between the ion species.
+A finite wavelength approximation was used for the screening cloud, with an effective Coulomb potential describing the electron-ion interactions.
+Form factors were calculated using the screened hydrogenic approximation Pauling and Sherman (1932).
+The partial static structure factors were calculated using a Debye-Hückel effective potential Gericke et al.
+(2010) within the hypernetted-chain approximation Wünsch et al.
+(2008).
+Appendix G Static structure factors
+In Fig.
+4, we compose the full Rayleigh weight at $T=150\,$eV and $\rho=20\,$g/cc into its individual components, i.e., the ionic static structure factor $S_{II}(q)$ [left] and the electron–ion static structure factor $S_{eI}(q)$ [right].
+Interestingly, we find overall very good agreement between all depicted models in the former, while there are more substantial disagreements in the latter.
+We thus conclude that the accurate description of the electronic localization around the ions poses the main challenge for the accurate description of $W_{\textnormal{R}}(q)$, whereas ionic correlations are less sensitive to the included microphysics.
+Finally, we show DFT-MD results for $S_{II}(q)$ at $T=150\,$eV and $\rho=5\,$g/cc for $N_{Be}=8$ (black) and $N_{Be}=20$ (red) beryllium atoms in Fig.
+5.
+Evidently, there are no finite-size effects except for the different $q$-grids, which is a consequence of momentum quantization in the finize simulation cell.